Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer.
Part 1: Using complete sentences, explain what Nakim did wrong.
Part 2: Show all your work to simplify the expression.
(You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). )

Answer by Theo(13342)

original expression is:
3*sqrt(2x) + x*sqrt(8x) - 5*sqrt(18x becomes:
3*sqrt(2x) + 2x*sqrt(2x) - 15*sqrt(2x).
you can factor out the sqrt(2x) to get:
sqrt(2x) * (3 + 2x - 15) which results in:
sqrt(2x) * (-12 + 2x)
the answer given was sqrt(2x) * (-10x)
looks like nakim combined 3 + 2x - 15 to get -10x.
the 3 and the -15 could be combined because they're like terms.
the 2x could not be combined because it was different from them. it included an x in the term which the others did not have.
a + a = 2a because they are like terms.
a + b = a + b because they are not like terms.